Jose has a board that is 44 inches long. He wishes to cut it into two pieces so that one piece will be 6 inches longer than the other. How long should the shorter piece be?
Answer by ikleyn(52781)

.
Jose has a board that is 44 inches long.
He wishes to cut it into two pieces so that one piece will be 6 inches longer than the other.
How long should the shorter piece be?
~~~~~~~~~~~~~~~~~

Let x be the length of one piece (of the shorter piece).

Then the length of the longer piece is (x+6), according to the problem.


Next, we write equation for the total length

    x + (x+6) = 44  inches.


Simplify and find x

    2x = 44 - 6

    2x = 38

     x = 38/2 = 19.


ANSWER.  The shorter piece is 19 inches long.  The longer piece is 19+6 = 25 inches long.


CHECK.   19 + 25 = 44 inches,  total.   ! Correct !

Solved.